# XPG-RL: Reinforcement Learning with Explainable Priority Guidance for Efficiency-Boosted Mechanical Search

### v1

@@ Abstract
Mechanical search (MS) in cluttered environments remains a significant challenge for autonomous manipulators, requiring long-horizon planning and robust state estimation under occlusions and partial observability. In this work, we introduce XPG-RL, a reinforcement learning framework that enables agents to efficiently perform MS tasks through explainable, priority-guided decision-making based on raw sensory inputs. XPG-RL integrates a task-driven action prioritization mechanism with a learned context-aware switching strategy that dynamically selects from a discrete set of action primitives such as target grasping, occlusion removal, and viewpoint adjustment. Within this strategy, a policy is optimized to output adaptive threshold values that govern the discrete selection among action primitives. The perception module fuses RGB-D inputs with semantic and geometric features to produce a structured scene representation for downstream decision-making. Extensive experiments in both simulation and real-world settings demonstrate that XPG-RL consistently outperforms baseline methods in task success rates and motion efficiency, achieving up to 4.5$\times$ higher efficiency in long-horizon tasks. These results underscore the benefits of integrating domain knowledge with learnable decision-making policies for robust and efficient robotic manipulation.
Keywords: Reinforcement Learning, Mechanical Search, Priority-Guided Decision-making
Figure 1: We propose XPG-RL, a reinforcement learning framework for mechanical search tasks. XPG-RL leverages task-guided action prioritization and learns context-aware switching over action primitives, effectively reducing redundant manipulations and improving task efficiency. The figure shows the manipulator successfully grasping a target object (banana) in a densely cluttered real-world scene.
1 Introduction
@@
Given the dynamic and unpredictable nature of MS tasks, predefining all possible scenarios is impractical. Learning-based approaches are essential for allowing robots to adapt to new situations and generalize beyond predefined rules [6]. Reinforcement learning (RL) enables robots to learn policies directly from visual inputs through trial-and-error exploration, without requiring explicit object pose knowledge, and is widely recognized as an effective approach for tackling MS tasks [7, 8, 9, 10]. However, conventional RL methods treat each action primitive independently via separate neural networks, neglecting action interdependencies [11, 12]. This limitation leads to cascading errors and can misguide decision-making in scenarios requiring long-horizon planning. Zenkri et al. observed that when the number of objects increased from 10 to 20, the efficiency of traditional RL methods dropped by 46.6% [13]. These challenges highlight the need for advancements in RL to enhance long-term MS performance and improve decision-making efficiency.
@@ 3.1 Perception
3.2 RL-based Decision-making
@@
$r_{t}=R(s_{t},a_{t},s_{t+1})=\begin{cases}1000,&\text{if the target object is successfully extracted;}\\ -100,&\text{if an infeasible action is selected;}\\ -1,&\text{otherwise.}\end{cases}$ (1)
@@ 4 Experiments
4.1 Simulation Experiments
4.1.1 Comparison to Baselines
We evaluate the performance of XPG-RL against several baseline methods across a range of scene complexities, defined by the number of objects in the environment: from 5 (simple cases) to 20 (highly cluttered cases). As object count increases, the scene becomes more occluded and challenging. Table 1 summarizes the task success rate and average motion number for each method.
Table 1: Success rate and efficiency across scenes. XPG-RL consistently achieves higher success rates with fewer actions than baselines, demonstrating robust and efficient performance.
@@ 4.1.1 Comparison to Baselines
To further investigate whether XPG-RL retains its advantage in short-horizon tasks, we conduct an additional evaluation on scenes with 3 to 5 objects. Table 2 presents the results. While both XPG-RL and MPGNet perform reasonably well in terms of task success, XPG-RL completes tasks with significantly fewer motions. This efficiency gain further underscores the benefit of learning-based high-level control, which enables XPG-RL to adapt its behavior based on task complexity and avoid redundant actions when a simple solution suffices.
@@
Table 2: Efficiency in short-horizon tasks. XPG-RL maintains higher success rates while significantly reducing the number of actions compared to MPGNet in simpler scenes, demonstrating adaptability and avoidance of redundant motions.
Method Task success rate [%] Average motion number
3 objects 4 objects 5 objects 3 objects 4 objects 5 objects
MPGNet 80 79 76 3.02 3.51 3.84
XPG-RL 92 86 81 1.70 2.12 2.31
4.1.2 Ablation Study
The results, summarized in Table 3, highlight the contribution of each module. Removing the priority guidance leads to a clear and consistent drop in both task success rate and motion efficiency, especially as scene complexity increases. This demonstrates that the structured action prioritization significantly improves decision quality, particularly in longer-horizon tasks. When comparing the full XPG-RL model to the No NBV Motion variant, we observe that NBV selection plays a relatively minor role in simpler scenes (e.g., 5 objects) but becomes increasingly important in complex and highly occluded environments. In such cases, NBV helps the agent improve visibility and locate graspable views of the target object, resulting in higher success rates and fewer failed attempts. Overall, the ablation study confirms that both priority guidance and NBV selection are essential for achieving robust, efficient performance, with their contributions becoming more pronounced as task complexity increases.
Table 3: Contribution of priority guidance and NBV. Removing priority guidance or NBV leads to lower success rates and efficiency, especially as scene complexity increases. This confirms that both components are critical for robust performance in cluttered environments.
@@ 4.1.2 Ablation Study
4.2 Real-World Experiments
The results are summarized in Table 4. Across all scenes, XPG-RL consistently achieves higher task success rates and requires fewer motions compared to baseline methods. In Scene 3, where the target object is fully occluded, XPG-RL effectively leverages NBV motions to reposition the camera, enabling successful identification and grasping. Rather than disturbing surrounding objects and worsening occlusion, it strategically selects a side viewpoint to minimize risk and maintain task feasibility. In Scene 4, the target object (banana) is partially occluded by a sugar box, where naive pushing or failed grasps could destabilize the scene and obscure the target further. XPG-RL avoids such failures through priority-guided decision-making that defers risky actions when predicted success probabilities fall below learned thresholds. This mechanism enables the agent to prioritize safer alternatives, such as NBV motions, and supports more informed, interpretable decisions. By learning context-aware thresholds, XPG-RL dynamically balances risk and reward based on the scene context, demonstrating strong generalization and reliable performance in real-world cluttered environments.
Table 4: Real-world performance across cluttered scenes. XPG-RL consistently achieves higher task success rates with fewer actions than baselines across four tabletop environments.
@@ 4.2 Real-World Experiments
5 Conclusion
@@
We propose XPG-RL, a reinforcement learning framework that enhances efficiency and robustness in mechanical search through structured action prioritization and adaptive action switching. XPG-RL consistently outperformed state-of-the-art methods in both simulation and real-world environments. In simulated cluttered scenes, XPG-RL achieved a 76% success rate with 10 objects and a 64% success rate with 20 objects. It also demonstrated up to a 4.5$\times$ improvement in relative efficiency for long-horizon tasks. In real-world experiments involving fully occluded scenes, XPG-RL completed all five trials with fewer actions, whereas state-of-the-art methods frequently failed. These results underscore the benefits of integrating domain knowledge with learnable decision-making policies for robust and efficient robotic manipulation.
6 Limitations
Despite its effectiveness, XPG-RL presents opportunities for further enhancement. First, the current priority structure and action set are carefully designed based on domain knowledge, ensuring reliable performance within targeted manipulation domains. While this structured approach contributes to sample efficiency, it may limit generalization to significantly different domains. Future work will explore leveraging large language models (e.g., GPT-4 [37] or LLaMA [38]) to automatically adapt or generate action primitives, further enhancing flexibility across domains. Second, XPG-RL selects the next-best-view (NBV) by predicting grasp quality across candidate viewpoints—a process that, while effective, can be computationally intensive. Future efforts aim to extend the framework toward jointly optimizing action selection and NBV strategies in an end-to-end fashion, which would reduce computational overhead and improve coordination between decision stages, enabling more scalable and adaptable performance across diverse task settings.
Acknowledgments
The authors would like to thank Dr. Katherine A. Skinner and Prof. Ram Vasudevan for their valuable feedback and support throughout this work.
Appendix A Training Details
A.1 Policy Training Hyperparameters
The RL agent is trained for a total of 10,000 steps using the PPO algorithm. The policy is optimized with a learning rate of $3\times 10^{-4}$, a discount factor ($\gamma$) of 0.99, and a generalized advantage estimation (GAE) parameter ($\lambda$) of 0.95. The training process uses a batch size of 64 and a minibatch size of 32, with each PPO update performed over 4 epochs. Clipping for the policy objective is set to 0.2 to ensure stable updates. Reward normalization and observation normalization are applied throughout training to enhance learning stability.
A.2 Network Architectures
The object recognition component of the perception pipeline begins with object detection using the Segment Anything Model (SAM) [30], which generates instance-level masks for each object in the scene. These masked images are then processed by PoseCNN [31], which utilizes a 13-layer VGG16-style convolutional backbone for feature extraction. This is followed by a semantic segmentation head with convolutional layers for per-pixel classification.
For policy learning, the PPO agent employs separate actor and critic networks, both implemented as multi-layer perceptrons (MLPs). The observation space consists of images with dimensions $2\times 640\times 480$ (target mask and object-level depth map, two frames stacked along the channel dimension). These inputs are first processed through a sequence of convolutional layers for feature extraction: a 32-filter convolutional layer with an $8\times 8$ kernel and stride 4, followed by a 64-filter layer with a $4\times 4$ kernel and stride 2, and a final 64-filter layer with a $3\times 3$ kernel and stride 1, each followed by $\mathrm{ReLU}$ activations. The resulting feature map is flattened and passed to the actor and critic MLPs, each comprising two hidden layers with 64 nodes per layer and $\tanh$ activation functions. The actor and critic maintain independent parameter sets, enabling separate learning of the policy and value functions. Orthogonal weight initialization is applied throughout the networks, with biases initialized to zero.
Appendix B Next-best-view Strategy
To determine the next-best-view (NBV), our method first constructs a Truncated Signed Distance Function (TSDF) representation of the current scene, which integrates depth information from previous observations into a 3D volumetric map. This TSDF map enables the system to simulate various candidate camera viewpoints without physically moving the robot.
For each candidate viewpoint, the method predicts the grasp quality score for the target object. After evaluating all candidates, the viewpoint with the highest predicted grasp quality is selected as the next camera pose.
This NBV strategy, inspired by affordance-based NBV approaches [39], enhances perception by increasing viewpoint diversity, leading to a more comprehensive understanding of cluttered scenes and improving generalization across tasks. Unlike traditional NBV methods that may trigger viewpoint adjustments at fixed intervals, our approach selectively activates NBV only when no higher-priority actions (e.g., grasping or occlusion removal) are viable. This conditional triggering avoids unnecessary motions and ensures that NBV is used specifically to gather informative observations, enhancing task efficiency without compromising performance.

### v2

@@ Abstract
Mechanical search (MS) in cluttered environments remains a significant challenge for autonomous manipulators, requiring long-horizon planning and robust state estimation under occlusions and partial observability. In this work, we introduce XPG-RL, a reinforcement learning framework that enables agents to efficiently perform MS tasks through explainable, priority-guided decision-making based on raw sensory inputs. XPG-RL integrates a task-driven action prioritization mechanism with a learned context-aware switching strategy that dynamically selects from a discrete set of action primitives such as target grasping, occlusion removal, and viewpoint adjustment. Within this strategy, a policy is optimized to output adaptive threshold values that govern the discrete selection among action primitives. The perception module fuses RGB-D inputs with semantic and geometric features to produce a structured scene representation for downstream decision-making. Extensive experiments in both simulation and real-world settings demonstrate that XPG-RL consistently outperforms baseline methods in task success rates and motion efficiency, achieving up to 4.5$\times$ higher efficiency in long-horizon tasks. These results underscore the benefits of integrating domain knowledge with learnable decision-making policies for robust and efficient robotic manipulation. The project page for XPG-RL is https://yitingzhang1997.github.io/xpgrl/.
footnotetext: Corresponding author. Email: yitzhang@umich.edu.
I Introduction
@@
Given the dynamic and unpredictable nature of MS tasks, predefining all possible scenarios is impractical. Learning-based approaches are essential for allowing robots to adapt to new situations and generalize beyond predefined rules [2, 38]. Reinforcement learning (RL) enables robots to learn policies directly from visual inputs through trial-and-error exploration, without requiring explicit object pose knowledge, and is widely recognized as an effective approach for tackling MS tasks [34, 12, 18, 29]. However, conventional RL methods treat each action primitive independently via separate neural networks, neglecting action interdependencies [7, 28]. This limitation leads to cascading errors and can misguide decision-making in scenarios requiring long-horizon planning. Zenkri et al. observed that when the number of objects increased from 10 to 20, the efficiency of traditional RL methods dropped by 46.6% [36]. These challenges highlight the need for advancements in RL to enhance long-term MS performance and improve decision-making efficiency.
@@ III-A Perception
III-B RL-based Decision-making
@@
$r_{t}=\begin{cases}1000,&\text{if the target is successfully extracted;}\\ -100,&\text{if an infeasible action is selected;}\\ -1,&\text{otherwise.}\end{cases}$ (1)
@@ IV Experiments
TABLE I: Success rate and efficiency across scenes. XPG-RL consistently achieves higher success rates with fewer actions than baselines, demonstrating robust and efficient performance.
@@
IV-A Simulation Experiments
IV-A1 Comparison to Baselines
We evaluate the performance of XPG-RL against several baseline methods across a range of scene complexities, defined by the number of objects in the environment: from 5 (simple cases) to 20 (highly cluttered cases). As object count increases, the scene becomes more occluded and challenging. Table I summarizes the task success rate and average motion number for each method.
@@ IV-A1 Comparison to Baselines
To further investigate whether XPG-RL retains its advantage in short-horizon tasks, we conduct an additional evaluation on scenes with 3 to 5 objects. The results are presented in Table IV in the appendix. While both XPG-RL and MPGNet perform reasonably well in terms of task success, XPG-RL completes tasks with significantly fewer motions. This efficiency gain further underscores the benefit of learning-based high-level control, which enables XPG-RL to adapt its behavior based on task complexity and avoid redundant actions when a simple solution suffices.
@@
TABLE II: Contribution of priority guidance and NBV. Removing priority guidance or NBV leads to lower success rates and efficiency, especially as scene complexity increases. This confirms that both components are critical for robust performance in cluttered environments.
@@
TABLE III: Real-world performance across cluttered scenes. XPG-RL consistently achieves higher task success rates with fewer actions than baselines across four tabletop environments.
@@
IV-A2 Ablation Study
The results, summarized in Table II, highlight the contribution of each module. Removing the priority guidance leads to a clear and consistent drop in both task success rate and motion efficiency, especially as scene complexity increases. This demonstrates that the structured action prioritization significantly improves decision quality, particularly in longer-horizon tasks. When comparing the full XPG-RL model to the No NBV Motion variant, we observe that NBV selection plays a relatively minor role in simpler scenes (e.g., 5 objects) but becomes increasingly important in complex and highly occluded environments. In such cases, NBV helps the agent improve visibility and locate graspable views of the target object, resulting in higher success rates and fewer failed attempts. Overall, the ablation study confirms that both priority guidance and NBV selection are essential for achieving robust, efficient performance, with their contributions becoming more pronounced as task complexity increases.
IV-B Real-World Experiments
The results are summarized in Table III. Across all scenes, XPG-RL consistently achieves higher task success rates and requires fewer motions compared to baseline methods. In Scene 3, where the target object is fully occluded, XPG-RL effectively leverages NBV motions to reposition the camera, enabling successful identification and grasping. Rather than disturbing surrounding objects and worsening occlusion, it strategically selects a side viewpoint to minimize risk and maintain task feasibility. In Scene 4, the target object (banana) is partially occluded by a sugar box, where naive pushing or failed grasps could destabilize the scene and obscure the target further. XPG-RL avoids such failures through priority-guided decision-making that defers risky actions when predicted success probabilities fall below learned thresholds. This mechanism enables the agent to prioritize safer alternatives, such as NBV motions, and supports more informed, interpretable decisions. By learning context-aware thresholds, XPG-RL dynamically balances risk and reward based on the scene context, demonstrating strong generalization and reliable performance in real-world cluttered environments.
V Conclusion
@@
We propose XPG-RL, a reinforcement learning framework that enhances efficiency and robustness in mechanical search through structured action prioritization and adaptive action switching. XPG-RL consistently outperformed state-of-the-art methods in both simulation and real-world environments. In simulated cluttered scenes, XPG-RL demonstrated up to a 4.5$\times$ improvement in relative efficiency. In real-world experiments involving fully occluded scenes, XPG-RL completed all five trials with fewer actions, whereas state-of-the-art methods frequently failed. These results underscore the benefits of integrating domain knowledge with learnable decision-making policies for robust and efficient robotic manipulation.